OpenStudy (anonymous):

What are all the roots of x^3 - 3x^2 - x +3 = 0?

OpenStudy (callisto):

1. do factorization x^2(x - 3) - (x -3) = 0 (x^2-1)(x-3) =0 (x-1)(x+1)(x-3) =0 2. for it to exist, either (x-1)=0 or (x+1)=0 or (x-3)=0 Can you solve them for here?

OpenStudy (callisto):

*delete the word ''for''

OpenStudy (anonymous):

The answers are -1 and 1, 1 and 3, -1,1, and 3, 1, -1, and 3. So the answer would be the last one correct?

OpenStudy (callisto):

-1,1, and 3, 1, -1, and 3. <- any differences?

OpenStudy (anonymous):

Yes, because the way you set it up the last one would only work because in the first one -1 - 1 is -2 not zero, so I guess it might be a trick question? I have no idea

OpenStudy (callisto):

*confused* I assume your 3rd choice is -1, 1 and 3 the 4th choice is 1, -1 and 3 Numerically, aren't they supposed to be the same?

OpenStudy (callisto):

And you need to know that you have to solve the 3 equations independently. (x-1)=0 or (x+1)=0 or (x-3)=0 x = 1 or x =-1 or x=3 So, you'll get 3 numbers.. and from your choices, I can hardly recognise their differences except the arrangement of the answers.

OpenStudy (anonymous):

Here is the question from the book.

OpenStudy (callisto):

lol you've made a typing mistake for the last choice :P

OpenStudy (anonymous):

Oh, lol sorry about that haha!

OpenStudy (callisto):

Can you get the answer now?

OpenStudy (anonymous):

Yes! Thanks for the help.

OpenStudy (callisto):

welcome :)